OpenStudy (anonymous):

What is the axis of symmetry for the graph of y=2x^2-6x+2? x=_________ ***No answer choices, and idk how to solve.. :( please explain?? thx :)

OpenStudy (anonymous):

\[x=-b/2a\]

OpenStudy (anonymous):

what do i do with that?

OpenStudy (anonymous):

a=2 b=-6 c=2

OpenStudy (anonymous):

so x= -(-6)/2(2) =6/4 =3/2 ???

OpenStudy (anonymous):

 is that right???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@jonask is that correct?? does x=3/2 ??

OpenStudy (anonymous):

this is like the x coordinate of the TP

OpenStudy (anonymous):

oh so i need to simplify 6/4 right?? and that is why x=3/2 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool! thanks :)